OpenStudy (ashleyisakitty):

A thundercloud has an electric charge of 43.2 C near the top of the cloud and -38.7 C near the bottom of the cloud. The magnitude of the electric force between these two charges is 3.95 x 106 N. What is the average separation between these charges? Please explain.

OpenStudy (nincompoop):

do you know the formula?

OpenStudy (ashleyisakitty):

not really.. ive studying coulombs but havent memorized the formulas within yet. F = (kq1q2)/r^2?

OpenStudy (nincompoop):

\[F _{electric}=k_c \frac{ q_1q_2 }{ r^2 }\]

OpenStudy (ashleyisakitty):

(kc=8.99×10^99N*m^2 / C^2)?

OpenStudy (nincompoop):

so in your case you have: \[k_c=\frac{ (43.2C)(-38.7C }{ r^2 }=3.95\times10^6N\]

OpenStudy (nincompoop):

get the value of r \[r=\sqrt{k_c \frac{ q_1q_2 }{ F _{electric} }}\]

OpenStudy (nincompoop):

ignore my = after the k_c

OpenStudy (nincompoop):

plug the values in @Ashleyisakitty and you have your answer :) = 1,951 m

OpenStudy (ashleyisakitty):

Awesome! Also, K=Q1Q2/D2 is forumla that works with Coulombs right?

OpenStudy (nincompoop):

tell me if you have the same

OpenStudy (ashleyisakitty):

Thanks Nin for helping out. I used the formula i listed and got the same answer. Even though formulas are just plugging in numbers theyre always a bitt confusing at first for me.